OpenStudy (anonymous):

A cone has a volume of 72 ft3. The volume of a cylinder with the same height and a congruent base is _____ ft3.

OpenStudy (diyadiya):

Volume of cone : \[\frac{1}{3}\pi r^2h=72ft\] Multiplying 3 on both sides \[\pi r^2h=216\] Volumes og cylinders = \(\pi r^2 h\) therefore 216

OpenStudy (diyadiya):

**Volume of cylinder=\(\pi r^2h\)